OpenStudy (jskinny0831):

Using the table, complete parts A, B, and C. Simple Interest Versus Compound Interest Simple Interest Compound Interest Year Price Year Price 0 $20,000.00 0 $20,000.00 10 $32,000.00 10 $35,816.95 20 A 20 B Part A: Find the third terms in the table, A and B, using a linear model (simple interest) and an exponential model (compound interest). Part B: Assuming it has the same annual interest rate rounded to the nearest cent, how much more is the investment worth using an exponential model versus a linear model?

OpenStudy (jskinny0831):

Part C: Discuss collaboratively: How would a rate of decay or decay factor affect the outcome in Parts A and B? Explain.

OpenStudy (jskinny0831):

i don't think they're gonna reply @CandyCove

OpenStudy (jskinny0831):

i just need to finish this to do good

OpenStudy (welshfella):

Simple interest linear equation slope = 32,000,000 - 20,000,000 ---------------------- 20 - 10 - what does that simplify to?

OpenStudy (jskinny0831):

1200000?

RhondaSommer (rhondasommer):



OpenStudy (welshfella):

yes so it can be modelled by the linear equation A = 1,200,000x + 20,000,000

RhondaSommer (rhondasommer):



OpenStudy (welshfella):

oh sorry I misread the quastion slope = 32,000.00 - 20,000.00 ------------------ 20 - 10 = 12,000.00 / 10 = 1,200.00

OpenStudy (jskinny0831):

so rhonda did it right?

OpenStudy (jskinny0831):

and okay so thats for part A?

OpenStudy (welshfella):

A = 1,200.00x + 20,000.00 to find the value of A plug in x = 20 and work it out

OpenStudy (welshfella):

sorry gotta go right now

OpenStudy (jskinny0831):

aw man so you only got partA

OpenStudy (jskinny0831):

@RhondaSommer is part A 44000?

RhondaSommer (rhondasommer):

yep

OpenStudy (jskinny0831):

really?

OpenStudy (jskinny0831):

its asking for the simple interest and the compound interest?

RhondaSommer (rhondasommer):

yep

OpenStudy (jskinny0831):

for the numbers of price A and B

RhondaSommer (rhondasommer):

letter B is compound letter A is simple

RhondaSommer (rhondasommer):

\(\Huge\cal\color{red}{Y}\color{#ff5C00}{o}\color{#FF8A00}{u}\)\(\Huge\cal\color{#FFB800}{~~d}\color{#FFE600}{i}\color{#E8FF00}{d}\)\(\Huge\cal\color{#D1FF00}{~~a}\color{#BAFF00}{m}\color{#00FF2E}{a}\color{#00FFA1}{z}\color{#00FFB8}{i}\color{#00FFE6}{n}\color{#00CFFF}{g}\)

OpenStudy (jskinny0831):

is A 1200.00 and B is 44000?

RhondaSommer (rhondasommer):

no a is 44000

OpenStudy (jskinny0831):

because i have to find the prices and I'm not sure

OpenStudy (jskinny0831):

ohh and what is B then

RhondaSommer (rhondasommer):

b is a seperate equation

RhondaSommer (rhondasommer):

I will not give you an immediate answer

OpenStudy (jskinny0831):

no thats not what I'm asking tho

RhondaSommer (rhondasommer):

$35,816.95 - $20,000.00 ____________________

OpenStudy (jskinny0831):

in the table its asking for Price A and B and 44000 is price A so I'm asking about price B

OpenStudy (jskinny0831):

15816.95?

OpenStudy (jskinny0831):

is it that? @RhondaSommer

OpenStudy (jskinny0831):

hello? @RhondaSommer

OpenStudy (jskinny0831):

@ganeshie8

OpenStudy (jskinny0831):

can you help?

RhondaSommer (rhondasommer):

add that to $35,816.95